# Neural Machine Unranking

### v2

@@ Abstract
We tackle the problem of machine unlearning within neural information retrieval, termed Neural Machine UnRanking (NuMuR) for short. Many of the mainstream task- or model-agnostic approaches for machine unlearning were designed for classification tasks. First, we demonstrate that these methods perform poorly on NuMuR tasks due to the unique challenges posed by neural information retrieval. Then, we develop a methodology for NuMuR named Contrastive and Consistent Loss (CoCoL), which effectively balances the objectives of data forgetting and model performance retention. Experimental results demonstrate that CoCoL facilitates more effective and controllable data removal than existing techniques.
Index Terms:
machine unlearning, neural ranking, information retrieval.
I Introduction
@@
It is therefore important to design methods for machine unlearning that can effectively deal with neural IR. Prominent existing model- and task-agnostic unlearning methods like Amnesiac Unlearning [9, 10] or Negative Gradient Removal [11, 12] (NegGrad) could be employed. However, these have been primarily designed for classification scenarios where it is typically possible to unlearn a class by deliberately damaging the model accuracy on the samples within that class; and Figure 1 illustrates that such unlearning strategies perform poorly in neural IR in the sense that reducing the performance of these models on the ‘forget set’ (i.e. on the data to be removed) incurs a severe performance loss on the ‘retain set’ (i.e. on the remaining data) and on test sets. We conjecture that this is due to strong dependencies in neural IR models, where removing individual data points disrupts learned patterns [13, 11, 14]. Another model-agnostic unlearning method is the teacher–student framework [15, 16] which was likewise originally designed for classification tasks. However, as we discuss in detail in Section II-C, a naïve application of this approach to neural IR fails because the relevance scores generated by neural ranking models cannot easily be normalized.
@@
    We formalise Neural Machine UnRanking (NuMuR) – the task of unlearning queries or documents within neural IR. We also provide two datasets to benchmark NuMuR.
@@
    We propose Contrastive and Consistent Loss (CoCoL) – a machine unlearning method specifically designed for the NuMuR task.
@@
    We demonstrate that CoCoL improves upon baseline methods in experimental validations. Specifically, CoCoL achieves controllable forgetting, enabling variable scales of data removal without markedly degrading the model’s performance across both retain and test sets.
II Background and problem definition
In this section, we provide, to our knowledge, the first formalisation of the task of machine unlearning within neural IR. We also explain why a naïve application of the teacher–student framework does not work in this context.
II-A Machine unlearning
Let $\mathcal{W}\subseteq\mathbb{R}^{d}$ the parameter space and let $\mathcal{S}$ be the universe of possible datasets. Let $M\colon\mathcal{S}\to\mathcal{W}$ be a learning algorithm which maps a training set $\text{{{S}}}\in\mathcal{S}$ to a model $w\in\mathcal{W}$. Learning algorithms may be random but we do not make this explicit in the notation. The trained model is then:
@@ II-A Machine unlearning
Informally, the teacher–student framework specifies the unlearning algorithm $U$ as using stochastic gradient-descent – initialised from $\mathcal{M}_{\text{{train}}}$ – to minimize (or at least decrease)
@@
II-B Unlearning in neural information retrieval
@@
A neural-ranking model $w\in\mathcal{W}$ is then trained to predict a relevance score $f_{w}(x)\in\mathbb{R}$ of some query–document pair $x=\langle q,d\rangle$. Relevance scores output by neural-ranking models are used to rank documents. Each document associated with a query is sorted by its score in (descending) order, so that higher scores correspond to a higher rank and thus earlier positions in the search results. Neural Machine UnRanking (NuMuR) is then the task that this model unlearns either queries or documents (or both):
@@
One of the difficulties encountered in NuMuR is that certain queries or documents may appear simultaneously in the retain set R and in the forget set F. For example, assume that the query: “The best one-week itinerary for a trip to London” is associated with two recommended itineraries (i.e., documents). If one itinerary’s owner recalls their answer, we must unlearn one query–document pair whilst maintaining the other.
@@
(a) Query removal. Here, queries $Q^{\prime}=\{q_{1},q_{2}\}$ are to be unlearned. Thus, $\text{{{F}}}=\{(\langle q_{1},d_{1}\rangle,y_{1,1}),(\langle q_{1},d_{2}\rangle,y_{1,2}),(\langle q_{2},d_{2}\rangle,y_{2,2})\}$, $\text{{{E}}}=\{(\langle q_{3},d_{2}\rangle,y_{2,2}),(\langle q_{4},d_{3}\rangle,y_{4,3})\}$ and $\text{{{D}}}=\{(\langle q_{4},d_{4}\rangle,y_{4,4}),(\langle q_{5},d_{5}\rangle,y_{5,5})\}$.
(b) Document removal. Here, documents $D^{\prime}=\{d_{2},d_{3}\}$ are to be unlearned. Thus, $\text{{{F}}}=\{(\langle q_{1},d_{2}\rangle,y_{1,2}),(\langle q_{2},d_{3}\rangle,y_{2,3}),(\langle q_{3},d_{2}\rangle,y_{3,2})\},(\langle q_{4},d_{3}\rangle,y_{4,3})\}$, $\text{{{E}}}=\{(\langle q_{1},d_{1}\rangle,y_{1,1}),(\langle q_{4},d_{4}\rangle,y_{4,4})\}$ and $\text{{{D}}}=\{(\langle q_{5},d_{5}\rangle,y_{5,5})\}$.
Figure 2: Illustration of machine unranking for the dataset $\text{{{S}}}=\{(\langle q_{1},d_{1}\rangle,y_{1,1}),(\langle q_{1},d_{2}\rangle,y_{1,2}),(\langle q_{2},d_{2}\rangle,y_{2,2}),(\langle q_{3},d_{2}\rangle,y_{2,2}),\allowbreak(\langle q_{4},d_{3}\rangle,y_{4,3}),(\langle q_{4},d_{4}\rangle,y_{4,4}),(\langle q_{5},d_{5}\rangle,y_{5,5})\}$.
II-C Breakdown of existing knowledge-distillation based unlearning algorithms in neural information retrieval
The teacher–student approach from Chundawat et al. [15] (reviewed at the end of Section II-A) was primarily designed for classification models where the objectives $L_{\mathcal{M},\mathsf{A}}(w)$ in (4) penalise the difference between class probabilities predicted by the student model $w$ and by the reference (‘teacher’) model $\mathcal{M}$. The teacher–student approach thus exploits the fact that the outputs of neural classification models are class probabilities which are always normalized, so that forgetting (e.g., of a class) can always be ensured by simply lowering the associated class probabilities. Unfortunately, the relevance scores generated by neural ranking models cannot typically be normalized so that the ranking implications of modifying relevance-score distributions are unclear (see, e.g., Figure 8 in Appendix Relevance score interval comparison); and this, as well as the fact that classification tasks do not involve entangled sets, causes the teacher–student approach to break down in neural IR. More specifically, we identify the following problems:
@@
    $\mathcal{M}_{\text{{init}}}$ cannot serve as ‘incompetent’ teacher. Due to the lack of normalization, the relevance scores on F are not necessarily lower under the ‘incompetent’ teacher model (e.g., $\mathcal{M}_{\text{{init}}}$) than under the trained model. For example, a relevance score of 30, say, might imply a high rank under the ‘competent’ teacher model, but a low rank under the incompetent teacher model (see Figure 8 in Appendix Relevance score interval comparison). Therefore, using $\mathcal{M}_{\text{{init}}}$ as the teacher for the forget set while employing $\mathcal{M}_{\text{{train}}}$ as a ‘competent’ teacher for the retain set may be counterproductive.
@@
    Controllable forgetting is challenging in neural ranking. In a $k$-class classification model, forgetting a specific sample is achieved by adjusting the model output (i.e., class probabilities $\in\mathbb{R}^{k}$) so that the probability of the correct class falls below $1/k$. This type of forgetting can be quantitatively assessed using the Kullback–Leibler divergence from the outputs of the teacher model to the student model [21, 16, 22]. However, in neural ranking models, achieving forgetting by manipulating the model output (i.e., relevance score $\in\mathbb{R}$, an unnormalized scalar) is challenging, due to the absence of a clear threshold or benchmark for adjusting these scores.
@@
    A naïve application of the teacher–student framework overlooks the entangled set. A challenge in NuMuR is that some queries or documents can appear in both the retain and the forget set, as formalised by the entangled set. Conventional teacher–student frameworks implement distinct strategies for the forget and retain sets, as summarized in (4), without accounting for the entangled set. However, effectively decoupling the learned relevance estimation patterns between the forget set and the entangled set using such unlearning methods is problematic. This issue will be evidenced in Figure 6, where we illustrate that teacher–student approach that ignores the entangled set yields inferior performance on both the retain and test sets.
III Proposed neural machine unranking methodology
In this section, we propose a new teacher–student framework for NuMuR, called Contrastive and Consistent Loss (CoCoL). To address the three challenges discussed at the end of Section II, CoCoL introduces the following elements.
@@ III Proposed neural machine unranking methodology
    To overcome the problem of using an ‘incompetent’ teacher model (such as $\mathcal{M}_{\text{{init}}}$) in the presence of unnormalized relevance scores, we attempt to reduce the relevance scores on the forget set (relative to the trained model, $\mathcal{M}_{\text{{train}}}$) whilst seeking to maintain the relevance scores on the retain set.
@@
    Given that relevance scores are not normalized, to enable controllable forgetting, we stop the unlearning iterations when
@@ III-A Objective
were $L_{\mathcal{M},\mathsf{A}}(w)$ is again some objective which penalises the discrepancy between $w$ and some reference model $\mathcal{M}$ on some dataset $\mathsf{A}$. Note that this objective differs from the standard teacher–student framework (4) in that the entangled set is moved into the first component. Note also that we say ‘decrease’ rather than ‘minimize’ because the unlearning is simply stopped when a pre-defined level of forgetting has been achieved (see below for details).
@@
$\Delta_{w,\mathcal{M}}^{\alpha,\beta}(x)=-\frac{\alpha f_{\mathcal{M}}(x)-f_{w}(x)+\beta}{f_{\mathcal{M}}(x)+f_{w}(x)},$ (8)
measures the discrepancy between the relevance score $f_{\mathcal{M}}(x)$ returned by some fixed reference (‘teacher’) model $\mathcal{M}$ and the relevance score $f_{w}(x)$ returned by the ‘student’ model $w$. Specifically, note that (8) decreases if the relevance score of the teacher model $\mathcal{M}$ is much higher than that of the student model $w$. In (8), $\alpha>0$ and $\beta\geq 0$ are tuning parameters whose choice will be discussed in Section III-B.
@@
    $\mathrm{ReLu}(\Delta_{w,\mathcal{M}_{\text{{train}}}}^{\alpha,\beta}(x))+\lvert\Delta_{w,\mathcal{M}_{\text{{train}}}}^{1,0}(x^{\prime})\rvert.$ (9)
@@
    $\lvert\Delta_{w,\mathcal{M}_{\text{{train}}}}^{1,0}(x^{+})\rvert+\lvert\Delta_{w,\mathcal{M}_{\text{{train}}}}^{1,0}(x^{-})\rvert.$ (10)
@@
III-B Choice of tuning parameters
[…]
III-C Stopping criteria of unranking
With appropriate settings of $\alpha$ and $\beta$ in (9), alternating between (9) and (10) ensures stable performance on both entangled and disjoint datasets, while performance on the forget set will progressively decline. Therefore, the optimal time to stop unlearning is when the performance on the forget set as measured by (6) reaches the pre-specified level $\delta>0$.
IV Experiments
IV-A Datasets
Currently, there are no existing IR datasets specifically designed for machine unlearning research. To address this, we propose curating datasets derived from established benchmark IR datasets. In selecting the appropriate datasets for NuMuR, our selection criteria focused on datasets that feature extensive one-to-many relevant query–document and document–query pairings, essential for evaluating NuMuR methodologies. An in-depth review of resources listed on ir-datasets.com identified two sources that fulfil these requirements: MS MARCO [18] and TREC Complex Answer Retrieval (TREC CAR) [25]. These sources were selected due to their large sample sizes and the presence of overlapping queries and documents. The sample ratio of the forget set, entangled set, and disjoint set is approximately $1:1:2$. Table I summarises the datasets.
TABLE I: Datasets created for this study.

### v3

@@ Abstract
We address the problem of machine unlearning in neural information retrieval (IR), introducing a novel task termed Neural Machine UnRanking (NuMuR). This problem is motivated by growing demands for data privacy compliance and selective information removal in neural IR systems. Existing task- or model- agnostic unlearning approaches, primarily designed for classification tasks, are suboptimal for NuMuR due to two core challenges: (1) neural rankers output unnormalised relevance scores rather than probability distributions, limiting the effectiveness of traditional teacher–student distillation frameworks; and (2) entangled data scenarios, where queries and documents appear simultaneously across both forget and retain sets, may degrade retention performance in existing methods. To address these issues, we propose Contrastive and Consistent Loss (CoCoL), a dual-objective framework. CoCoL comprises (1) a contrastive loss that reduces relevance scores on forget sets while maintaining performance on entangled samples, and (2) a consistent loss that preserves accuracy on retain set. Extensive experiments on MS MARCO and TREC CAR datasets, across four neural IR models, demonstrate that CoCoL achieves substantial forgetting with minimal retain and generalisation performance loss. Our method facilitates more effective and controllable data removal than existing techniques.
Keywords machine unlearning  $\cdot$ data removal  $\cdot$ neural ranking  $\cdot$ information retrieval  $\cdot$ controllable forgetting.
1 Introduction
@@
It is therefore important to design methods for machine unlearning that can effectively deal with neural IR. Prominent existing model- and task- agnostic unlearning methods like Amnesiac Unlearning [9, 10] or Negative Gradient Removal [11, 12] (NegGrad) could be employed. However, these have been primarily designed for classification scenarios where it is typically possible to unlearn a class by deliberately damaging the model accuracy on the samples within that class; and Figure 1 illustrates that such unlearning strategies perform poorly in neural IR in the sense that reducing the performance of these models on the ‘forget set’ (i.e. on the data to be removed) incurs a severe performance loss on the ‘retain set’ (i.e. on the remaining data) and on test sets. We conjecture that this is due to strong dependencies in neural IR models, where removing individual data points disrupts learned patterns [13, 11, 14]. Another model-agnostic unlearning method is the teacher–student framework [15, 16] which was likewise originally designed for classification tasks. However, as we discuss in detail in Section 2.3, a naïve application of this approach to neural IR fails because the relevance scores generated by neural ranking models cannot easily be normalised.
@@
    Formalisation of Neural Machine UnRanking (NuMuR) – the task of unlearning queries or documents within neural IR, along with two datasets to benchmark NuMuR.
@@
    Introduction of Contrastive and Consistent Loss (CoCoL) – a machine unlearning method specifically designed for the NuMuR task.
@@
    Experimental validation demonstrating that CoCoL improves upon baseline methods. Specifically, CoCoL achieves controllable forgetting, enabling variable scales of data removal without markedly degrading the model’s performance across both retain and test sets.
2 Background and problem definition
In this section, we provide the first formal definition of machine unlearning within neural IR and explain why existing teacher–student frameworks are inadequate for this domain.
2.1 Machine unlearning
Let $\mathcal{W}\subseteq\mathbb{R}^{d}$ be the parameter space and let $\mathcal{S}$ be the universe of possible datasets. Let $M\colon\mathcal{S}\to\mathcal{W}$ be a learning algorithm which maps a training set $\text{{{S}}}\in\mathcal{S}$ to a model $w\in\mathcal{W}$. Learning algorithms may be random but we do not make this explicit in the notation. The trained model is then:
@@ 2.1 Machine unlearning
Informally, the teacher–student framework specifies the unlearning algorithm $U$ as using stochastic gradient-descent – initialised from $\mathcal{M}_{\text{{train}}}$ – to minimise (or at least decrease)
@@
2.2 Unlearning in neural information retrieval
@@
A neural-ranking model $w\in\mathcal{W}$ is then trained to predict a relevance score $f_{w}(x)\in\mathbb{R}$ of some query–document pair $x=\langle q,d\rangle$. Relevance scores output by neural-ranking models are used to rank documents. Each document associated with a query is sorted by its score in (descending) order so that higher scores correspond to a higher rank and thus earlier positions in the search results.
Neural Machine UnRanking (NuMuR) is then the task that this model unlearns either queries or documents (or both):
@@
One of the difficulties encountered in NuMuR is that certain queries or documents may appear simultaneously in the retain set R and in the forget set F. For example, assume that the query: “The best one-week itinerary for a trip to London” is associated with two recommended itineraries (i.e., documents). Consider if one itinerary’s owner recalls their answer, we must unlearn one query–document pair whilst maintaining the other.
@@
(a) Query removal.
(b) Document removal.
Figure 2: Illustration of machine unranking for the dataset. $\text{{{S}}}=\{(\langle q_{1},d_{1}\rangle,y_{1,1}),(\langle q_{1},d_{2}\rangle,y_{1,2}),(\langle q_{2},d_{2}\rangle,y_{2,2}),(\langle q_{3},d_{2}\rangle,y_{2,2}),\allowbreak(\langle q_{4},d_{3}\rangle,y_{4,3}),(\langle q_{4},d_{4}\rangle,y_{4,4}),(\langle q_{5},d_{5}\rangle,y_{5,5})\}$.   (a) Query removal. Here, queries $Q^{\prime}=\{q_{1},q_{2}\}$ are to be unlearned. Thus, $\text{{{F}}}=\{(\langle q_{1},d_{1}\rangle,y_{1,1}),(\langle q_{1},d_{2}\rangle,y_{1,2}),(\langle q_{2},d_{2}\rangle,y_{2,2})\}$, $\text{{{E}}}=\{(\langle q_{3},d_{2}\rangle,y_{2,2}),(\langle q_{4},d_{3}\rangle,y_{4,3})\}$ and $\text{{{D}}}=\{(\langle q_{4},d_{4}\rangle,y_{4,4}),(\langle q_{5},d_{5}\rangle,y_{5,5})\}$.   (b) Document removal. Here, documents $D^{\prime}=\{d_{2},d_{3}\}$ are to be unlearned. Thus, $\text{{{F}}}=\{(\langle q_{1},d_{2}\rangle,y_{1,2}),(\langle q_{2},d_{3}\rangle,y_{2,3}),(\langle q_{3},d_{2}\rangle,y_{3,2}),(\langle q_{4},d_{3}\rangle,y_{4,3})\}$, $\text{{{E}}}=\{(\langle q_{1},d_{1}\rangle,y_{1,1}),(\langle q_{4},d_{4}\rangle,y_{4,4})\}$ and $\text{{{D}}}=\{(\langle q_{5},d_{5}\rangle,y_{5,5})\}$.
2.3 Breakdown of existing knowledge-distillation based unlearning algorithms in neural information retrieval
The teacher–student approach from Chundawat et al. [15] (reviewed in Section 2.1) was primarily designed for classification models where the objectives $L_{\mathcal{M},\mathsf{A}}(w)$ in (4) penalise the difference between class probabilities predicted by the student model $w$ and by the reference (‘teacher’) model $\mathcal{M}$. The teacher–student approach thus exploits the fact that the outputs of neural classification models are class probabilities which are always normalised, so that forgetting (e.g., of a class) can always be ensured by simply lowering the associated class probabilities. However, the relevance scores generated by neural ranking models cannot typically be normalised so that the ranking implications of modifying relevance-score distributions are unclear (see, e.g., Figure 8 in Appendix Relevance score interval comparison); and this, as well as the fact that classification tasks do not involve entangled sets, causes the teacher–student approach to break down in neural IR. More specifically, we identify the following problems:
@@
    $\mathcal{M}_{\text{{init}}}$ cannot serve as ‘incompetent’ teacher. Due to the lack of normalisation, the relevance scores on F are not necessarily lower under the ‘incompetent’ teacher model (e.g., $\mathcal{M}_{\text{{init}}}$) than under the trained model. For example, a relevance score of 30, say, might imply a high rank under the ‘competent’ teacher model, but a low rank under the incompetent teacher model (see Figure 8 in Appendix Relevance score interval comparison). Therefore, using $\mathcal{M}_{\text{{init}}}$ as the teacher for the forget set while employing $\mathcal{M}_{\text{{train}}}$ as a ‘competent’ teacher for the retain set may be counterproductive.
@@
    Controllable forgetting is challenging in neural ranking. In a $k$-class classification model, forgetting a specific sample is achieved by adjusting the model output (i.e., class probabilities $\in\mathbb{R}^{k}$) so that the probability of the correct class falls below $1/k$. This type of forgetting can be quantitatively assessed using the Kullback–Leibler divergence from the outputs of the teacher model to the student model [21, 16, 22]. However, in neural ranking models, achieving forgetting by manipulating the model output (i.e., relevance score $\in\mathbb{R}$, an unnormalised scalar) is challenging, due to the absence of a clear threshold or benchmark for adjusting these scores.
@@
    A naïve application of the teacher–student framework overlooks the entangled set. A challenge in NuMuR is that some queries or documents can appear in both the retain and the forget set, as formalised by the entangled set. Conventional teacher–student frameworks implement distinct strategies for the forget and retain sets, as summarised in (4), without accounting for the entangled set. However, effectively decoupling the learned relevance estimation patterns between the forget set and the entangled set using such unlearning methods is problematic. This issue will be evidenced in Figure 7, where we illustrate that teacher–student approach that ignores the entangled set yields inferior performance on both the retain and test sets.
3 Proposed neural machine unranking methodology
In this section, we propose a new teacher–student framework for NuMuR, called Contrastive and Consistent Loss (CoCoL). To address the three challenges discussed at the end of Section 2, CoCoL introduces the following elements.
@@ 3 Proposed neural machine unranking methodology
    To overcome the problem of using an ‘incompetent’ teacher model (such as $\mathcal{M}_{\text{{init}}}$) in the presence of unnormalised relevance scores, we attempt to reduce the relevance scores on the forget set (relative to the trained model, $\mathcal{M}_{\text{{train}}}$) whilst seeking to maintain the relevance scores on the retain set.
@@
    Given that relevance scores are not normalised, to enable controllable forgetting, we stop the unlearning iterations when
@@ 3.1 Objective
were $L_{\mathcal{M},\mathsf{A}}(w)$ is again some objective which penalises the discrepancy between $w$ and some reference model $\mathcal{M}$ on some dataset $\mathsf{A}$. Note that this objective differs from the standard teacher–student framework (4) in that the entangled set is moved into the first component. Note also that we say ‘decrease’ rather than ‘minimise’ because the unlearning is simply stopped when a pre-defined level of forgetting has been achieved (see below for details).
@@
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta_{w,\mathcal{M}}(x)=\frac{f_{\mathcal{M}}(x)-f_{w}(x)}{f_{\mathcal{M}}(x)+f_{w}(x)},}}$ (8)
measures the discrepancy between the relevance score $f_{\mathcal{M}}(x)$ returned by some fixed reference (‘teacher’) model $\mathcal{M}$ and the relevance score $f_{w}(x)$ returned by the ‘student’ model $w$. This measure supports maintaining relevance consistency for.
To facilitate targeted forgetting, for each $(x,y)=(\langle q,d\rangle,y)\in\text{{{F}}}$, we define the query-specific set of samples in S:
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{{{S}}}_{x}=\text{{{S}}}_{\langle q,d\rangle}\coloneqq\{\langle q^{*},d^{*}\rangle\in\mathcal{Q}\times\mathcal{D}\mid\exists\left(\langle q^{*},d^{*}\rangle,y^{*}\right)\in\text{{{S}}}}}:$ (9)
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q^{*}=q\}}}$ (10)
and compute the minimal relevance score for query $q$ in the teacher model:
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{\text{min}}_{\mathcal{M}}(x)=\min_{x\in\text{{{S}}}_{x}}f_{\mathcal{M}}(x)}}$ (11)
We then define an adjusted discrepancy measure to suppress the relevance score for forget set:
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta^{\text{min}}_{w,\mathcal{M}}(x)=-\frac{s^{\text{min}}_{\mathcal{M}}(x)-f_{w}(x)}{s^{\text{min}}_{\mathcal{M}}(x)+f_{w}(x)},}}$ (12)
Here, (12) prioritises reducing $f_{w}(x)$ toward the minimal relevance score $s^{\min}_{\mathcal{M}}(x)$ produced by $\mathcal{M}$ for query $q$.
@@
    ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{ReLu}(\Delta^{\text{min}}_{w,\mathcal{M}_{\text{{train}}}}(x))+\lvert\Delta_{w,\mathcal{M}_{\text{{train}}}}(x^{\prime})\rvert.}}$ (13)
@@
    ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lvert\Delta_{w,\mathcal{M}_{\text{{train}}}}(x^{+})\rvert+\lvert\Delta_{w,\mathcal{M}_{\text{{train}}}}(x^{-})\rvert.}}$ (14)
@@
3.2 Stopping criteria of unranking
Algorithm 1 NuMuR procedure based on CoCoL
1:Trained model $\mathcal{M}_{\text{{train}}}$, forget set F, entangled set E, disjoint set D, target forgetting level $\delta$, maximum epochs $N$
2:Updated model $w^{*}$
3:Set model parameters $w_{0}\leftarrow\mathcal{M}_{\text{{train}}}$
4:for $e=1$ to $N$ do
5:     // Phase 1: Forget and entangled set updates
6:     for all $(x,y)\in\text{{{F}}}\cup\text{{{E}}}$ do
7:         Compute loss $\mathcal{L}_{\text{forget}}(w_{e})$ using Equation (13)
8:         Update $w_{e}$ by gradient step on $\mathcal{L}_{\text{forget}}$
9:     end for
10:     // Phase 2: Disjoint set updates
11:     for all $(x,y)\in\text{{{D}}}$ do
12:         Compute loss $\mathcal{L}_{\text{retain}}(w_{e})$ using Equation (14)
13:         Update $w_{e}$ by gradient step on $\mathcal{L}_{\text{retain}}$
14:     end for
15:     Compute $\mathrm{MRR}_{\text{{{F}}}}(w_{e})$
16:     if $\mathrm{MRR}_{\text{{{F}}}}(w_{e})\leq\delta$ then
17:         break // Early stopping
18:     end if
19:end for
20:return Final model $w^{*}\leftarrow w_{e}$
As shown in algorithm 1, theocratically, alternating between (13) and (14) ensures stable performance on both entangled and disjoint datasets, while performance on the forget set will progressively decline. Therefore, the optimal time to stop unlearning is when the performance on the forget set as measured by (6) reaches the pre-specified level $\delta>0$.
4 Experiments
4.1 Datasets
Currently, there are no existing IR datasets specifically designed for machine unlearning research. To address this, we propose curating datasets derived from established benchmark IR datasets. In selecting the appropriate datasets for NuMuR, our selection criteria focused on datasets that feature extensive one-to-many relevant query–document and document–query pairings, essential for evaluating NuMuR methodologies. An in-depth review of resources listed on ir-datasets.com identified two sources that fulfil these requirements: MS MARCO [18] and TREC Complex Answer Retrieval (TREC CAR) [25]. These sources were selected due to their large sample sizes and the presence of overlapping queries and documents. Table 1 summarises the datasets.
Table 1: Datasets created for this study.